# Multi-Channel Hypergraph Contrastive Learning for Matrix Completion

### v1

@@ Abstract.
Rating is a typical user explicit feedback that visually reflects how much a user likes a related item. The (rating) matrix completion is essentially a rating prediction process, which is also a significant problem in recommender systems. Recently, graph neural networks (GNNs) have been widely used in matrix completion, which captures users’ preferences over items by formulating a rating matrix as a bipartite graph. However, existing methods are susceptible due to data sparsity and long-tail distribution in real-world scenarios. Moreover, the messaging mechanism of GNNs makes it difficult to capture high-order correlations and constraints between nodes, which are essentially useful in recommendation tasks. To tackle these challenges, we propose a Multi-Channel Hypergraph Contrastive Learning framework for matrix completion, named MHCL. Specifically, MHCL adaptively learns hypergraph structures to capture high-order correlations between nodes and jointly captures local and global collaborative relationships through attention-based cross-view aggregation. Additionally, to consider the magnitude and order information of ratings, we treat different rating subgraphs as different channels, encourage alignment between adjacent ratings, and further achieve the mutual enhancement between different ratings through multi-channel cross-rating contrastive learning. Extensive experiments on five public datasets demonstrate that the proposed method significantly outperforms the current state-of-the-art approaches.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Recently, the exponential growth of information has made matrix completion essential for e-commerce and social media platforms (Zhang et al., 2019; Ricci et al., 2010; Xia et al., 2020). These systems typically process two types of user feedback: explicit and implicit (Chen and Wang, 2022). Explicit feedback directly captures user preferences—for example, ratings on a matrix, where each entry quantifies how much a user likes a specific item. In contrast, implicit feedback indirectly reflects preferences, typically inferred as positive signals. This includes behavioral data such as click records, items added to shopping carts, and user favorites (Palomares et al., 2018; Raza and Ding, 2019). Matrix completion is a fundamental problem with various applications, ranging from missing traffic data imputation, protein relation prediction, image restoration, to recommender systems. Specifically, in recommender systems, the aim is to predict missing entries in a partially observed user-item matrix using available explicit or implicit feedback.
@@
Matrix factorization (Candes and Recht, 2012; Kalofolias et al., 2014; Xu et al., 2013) is a traditional approach to matrix completion that decomposes the user-item rating matrix into latent factors representing users and items. In recent years, GNNs have gained prominence in recommender systems (Monti et al., 2017; Shen et al., 2021; Zhang and Chen, 2019; Zhang et al., 2022). These methods model the rating matrix as a bipartite graph, reformulating the matrix completion problem as a link prediction task on this graph. Inductive matrix completion methods, such as IGMC (Zhang and Chen, 2019), utilize GNNs to learn local graph patterns from 1-hop subgraphs around target edges, enabling generalization to unseen nodes. Building on this idea, GIMC (Zhang et al., 2022) introduces edge embeddings and hyperbolic geometry to further enhance the modeling of bipartite graphs. Moreover, some review-based models (Wu et al., 2019; Gao et al., 2020a; Shuai et al., 2022) leverage auxiliary review information to enrich the learning of embeddings, complementing the structure-based methods.
Despite their remarkable success, existing GNN-based matrix completion methods face several key challenges. First, existing GNN-based methods struggle to capture high-order correlations and complex constraints due to their reliance on existing edges in bipartite graphs, which often require stacking multiple layers and can introduce noise, leading to over-smoothing issues (Min et al., 2020). Second, these methods are inherently vulnerable to data sparsity and long-tail distributions (Singh, 2020), where a few highly active nodes dominate interactions, resulting in popularity bias and diminished effectiveness for less active nodes. Many approaches also overlook the magnitude and ordinal nature of ratings. For example, a rating of ‘4’ is closer to ‘5’ than to ‘1’. Although some recent methods (Zhang and Chen, 2019; Shen et al., 2021) incorporate regularization to address certain limitations, they often fail to fully account for these aspects within the model design.
@@
Motivated by the need to address the aforementioned challenges, we propose a novel Multi-Channel Hypergraph Contrastive Learning (MHCL) framework for matrix completion. First, we partition the bipartite graph into multiple subgraphs based on different types of rating edges using a subgraph partition and graph convolution module. During training, an adaptive hypergraph structure learning module dynamically learns hypergraph structures for these subgraphs at different layers, capturing high-order correlations between nodes to overcome the first challenge. Second, we introduce a multi-channel cross-rating contrastive learning module that treats each rating subgraph as a distinct channel, promoting mutual enhancement by aligning adjacent views. Third, to address the data sparsity and long-tail distribution challenge, we design an attention-based cross-view fusion module that strengthens interactions between different rating channels. Finally, an attention-based cross-view aggregation module learns node representations collaboratively from both local and global perspectives, fusing them into the final representation. MHCL demonstrates significant performance improvements, achieving up to +9.04% lift in MSE over SOTA baselines for matrix completion and a +10.31% increase in NDCG@10 for recommendation tasks on the ML-100K dataset.
@@
  • MHCL leverages hypergraph learning to partition the user-item relationship graph into distinct subgraphs, representing various types of rating edges. It dynamically learns hypergraph structures to capture higher-order relationships during training.
  • MHCL introduces a multi-channel cross-rating contrastive learning, treating different rating subgraphs as unique perspectives. This approach promotes alignment between adjacent perspectives, enhancing the modeling of interrelationships between diverse rating types and improving prediction accuracy.
  • MHCL includes an attention-based cross-view aggregation module to integrate information from various rating subgraphs. It employs contrastive learning to collaboratively learn node representations from both local and global perspectives, ultimately fusing them to obtain the final node representations. This multi-view approach comprehensively captures node features, making contributions to improving recommendations.
  • We conduct comprehensive experiments to evaluate the performance of our MHCL model on matrix completion and recommendation tasks, along with diverse baselines, using multiple benchmark datasets.
@@ 2.1. GNNs for Matrix Completion
Recent advancements in graph neural networks (GNNs) have significantly enhanced matrix completion tasks. GNN-based methods interpret the matrix as a bipartite graph, where observed ratings or purchases are represented by edges, transforming matrix completion into a graph-based predictive task. Traditional GNN approaches (Kipf and Welling, 2016; Hamilton et al., 2017; Veličković et al., 2017) use multi-hop convolutions to capture high-order relations from user-item graphs. ST-GCN (Yan et al., 2018), NGCF (Wang et al., 2019) and PinSage (Ying et al., 2018) explore high-order connectivity through multiple graph convolutions, while LightGCN (He et al., 2020) emphasizes neighborhood aggregation by omitting non-linear transformations. Inductive methods like IDCF-NN (Wu et al., 2021b) and IDCF-GC (Wu et al., 2021b) utilize matrix factorization and neural information transfer for embedding computation. GC-MC (Berg et al., 2017) refines node representations using a GNN encoder and decoder, and IMC-GAE (Shen et al., 2021) improves generalization with role-aware attributes and one-hot encoding. IGMC (Zhang and Chen, 2019) extracts enclosing subgraphs to learn local patterns, while GIMC (Zhang et al., 2022) addresses data sparsity with hyperbolic regression. RMG (Wu et al., 2019) integrates graph signals with review information, and RGCL (Shuai et al., 2022) enhances this with semantic review features and contrastive learning (CL). MAGCL (Wang et al., 2023) further refines interaction modeling using multi-view representation, graph contrastive learning, and multi-task learning to capture fine-grained semantic information.
@@
In summary, these advancements within the realm of graph neural networks have ushered in transformative approaches to matrix completion, showcasing their potential to revolutionize recommender systems.
@@ 2.2. Hypergraph Learning for Recommendation
Inspired by the flexibility of hypergraphs to overcome the constraints of fixed graph structures in message passing (Gao et al., 2020b; Zhou et al., 2006; Bu et al., 2010; Yang et al., 2019), recent advancements in recommender systems have integrated hypergraph connections to improve relational learning. DualHGCN (Xue et al., 2021) converts multiplex bipartite networks into homogeneous hypergraphs and uses spectral hypergraph convolutional operators for effective intra- and inter-domain information exchange. Similarly, MHCN (Yu et al., 2021) utilizes social-based hypergraph convolutional networks to exploit high-order user relations in a multi-channel framework, while DHCF (Ji et al., 2020) focuses on modeling high-order correlations between users and items. However, these methods rely on manually crafted, static heuristic hypergraph structures.
To address this, recent approaches have explored dynamic hypergraph learning. For example, DHSL (Zhang et al., 2018) uses labels as spatial feature vectors for clustering, and DHGNN (Jiang et al., 2019) applies K-NN and K-means clustering to update hypergraph structures based on local and global features. Additionally, HCCF (Xia et al., 2022) and DHLCF (Wei et al., 2022) use collaborative signals and topological information to dynamically determine hyperedge assignments with MLPs. In contrast, our approach dynamically learns a hypergraph for each rating subgraph, capturing global collaborative relationships at a finer level among homogeneous nodes. Moreover, in addition to the hypergraph structure, DGCLR (Ren et al., 2022), VGCL (Yang et al., 2023) and ReHCL (Wang et al., 2024) consider fusing user comments and ratings for recommendation and introduce the user’s emotional tendency into the node representation learning to improve the recommendation performance.

### v2

@@ Abstract.
Rating is a typical user’s explicit feedback that visually reflects how much a user likes a related item. The (rating) matrix completion is essentially a rating prediction process, which is also a significant problem in recommender systems. Recently, graph neural networks (GNNs) have been widely used in matrix completion, which captures users’ preferences over items by formulating a rating matrix as a bipartite graph. However, existing methods are susceptible due to data sparsity and long-tail distribution in real-world scenarios. Moreover, the messaging mechanism of GNNs makes it difficult to capture high-order correlations and constraints between nodes, which are essentially useful in recommendation tasks. To tackle these challenges, we propose a Multi-Channel Hypergraph Contrastive Learning framework for matrix completion, named MHCL. Specifically, MHCL adaptively learns hypergraph structures to capture high-order correlations between nodes and jointly captures local and global collaborative relationships through attention-based cross-view aggregation. Additionally, to consider the magnitude and order information of ratings, we treat different rating subgraphs as different channels, encourage alignment between adjacent ratings, and further achieve the mutual enhancement between different ratings through multi-channel cross-rating contrastive learning. Extensive experiments on eight publicly available real-world datasets demonstrate that our proposed method significantly outperforms the current state-of-the-art approaches. The source code of our model is available at https://anonymous.4open.science/r/MHCL-25DC.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Mathematics of computing Graph algorithmsccs: Computing methodologies Learning latent representations
@@ 1. Introduction
In recent years, with the explosion of information, recommender systems have become an indispensable tool for e-commerce and social media platforms (Zhang et al., 2019; Ricci et al., 2010; Xia et al., 2020). In recommender systems, users’ feedback on items is divided into explicit and implicit feedback (Chen and Wang, 2022; Li et al., 2025). Implicit feedback means that the user does not explicitly like or dislike the item, which reveals that this type of data is generally considered positive feedback for all interactions (Palomares et al., 2018; Raza and Ding, 2019). For example, click records, shopping carts of shopping sites, and favorites or forwarding of items are part of the implicit feedback. Explicit feedback reflects how much a user likes a related item. In particular, given a partially observed matrix of users to items, the values in the matrix represent the users’ ratings of items. Then the process of completing the matrix can be regarded as the prediction process of ratings.
@@
As a kind of traditional matrix completion method, matrix factorization (Candes and Recht, 2012; Kalofolias et al., 2014; Xu et al., 2013) aims to decompose the user-item rating matrix into the user latent factors and item latent factors that profile users and items accurately. Recently, Graph Neural Networks (GNNs) have been widely used in recommender systems (Monti et al., 2017; Shen et al., 2021; Zhang and Chen, 2019; Zhang et al., 2022), these GNN-based methods formulate the rating matrix as a bipartite graph, and ratings in the matrix are treated as different types of edges, thus transforming the matrix completion problem into a link prediction (Martínez et al., 2016; Kumar et al., 2020; Arrar et al., 2024) problem on the bipartite graph. Researchers introduce inductive matrix completion methods, which can easily generalize to the nodes unseen during training and enhance the generalization of the model. IGMC (Zhang and Chen, 2019) extracts a 1-hop subgraph around all target edges, relabels nodes as the initial features in each subgraph, and applies GNN to each subgraph to learn the local graph pattern that can be generalized to the invisible graphs. GIMC (Zhang et al., 2022) introduces edge embeddings to model the semantic properties of different types of explicit feedback based on extracting subgraphs, and further improves the modeling ability of bipartite graphs by introducing hyperbolic geometry as the vertex-level embedding space. In addition, there are some review-based models (Wu et al., 2019; Gao et al., 2020a; Shuai et al., 2022) that utilize the auxiliary review information accompanied with user ratings to enrich the embedding learning.
Despite the remarkable success, existing GNN-based matrix completion methods still suffer from two significant challenges: First, existing GNN-based methods struggle to capture high-order correlations and complex constraints between nodes. GNN-based methods rely on existing edges in the bipartite graph to learn the node representations. Still, for high-order correlations and constraints, traditional GNNs need to stack multiple layers to capture, which not only brings much noise but makes it difficult to mitigate over-smoothing issues (Min et al., 2020). Second, user-item interaction data is usually sparse and follows long-tail and power-law distributions, while GNN-based methods are inherently vulnerable to data sparsity and long-tail distributions. The long-tail distribution problem in recommender systems (McNee et al., 2006) refers to the imbalance in the number of interactions between different nodes; that is, a few highly active nodes dominate interactions, resulting in popularity bias and diminished effectiveness for less active nodes. For rating recommendations, the long-tail distribution will also lead to a rating imbalance. Experimental results in (Mahadevan and Arock, 2021; Wei et al., 2021; Zhao et al., 2020) show that the imbalance distribution will cause popularity bias (Wu and Huang, 2022), which will lead to the model being more biased towards active nodes and ratings, and reduce the recommendation effect of inactive nodes and ratings. In Figure 1, we illustrate the long-tail distribution problem through Amazon dataset, which has the problem of the imbalance in the number of interactions and rating imbalance. We can see that the existing models (IMC-GAE (Shen et al., 2021) and GIMC (Zhang et al., 2022) as examples) have significantly higher prediction errors for tail classes (i.e., 1-score, 2-score, ‘Inactive’ group) than for head classes (i.e., 3-score, 4-score, 5-score, ‘Normal’ and ‘Active’ group), while our model can reduce the prediction error for tail classification. Unlike implicit feedback, rating information is based on the magnitude and order (rating ‘4’ is closer to ‘5’ rather than ‘1’), which may be ignored by many matrix completion methods. Although there are some recent methods (Zhang and Chen, 2019; Shen et al., 2021) that use regularization to mitigate these problems, they still do not consider these from a model perspective.
Figure 1. The imbalance in the number of interactions between different nodes and rating imbalance caused by the long-tail distribution problem. Specifically, we designate 80% of users with the fewest interactions as ‘Inactive’, classify the top 5% of users with the most interactions as ‘Active’, and assign the remaining users as ‘Normal’. Then we calculate the mean square error for each interaction group and rating group.
@@
Desired by addressing the challenges mentioned above, we propose a novel Multi-Channel Hypergraph Contrastive Learning framework for the matrix completion problem, named MHCL. First, we divide the bipartite graph into different subgraphs according to different types of rating edges through subgraph partition and graph convolution module, and then dynamically learn hypergraph structures for different subgraphs at different layers during the training process through an adaptive hypergraph structure learning module, to solve the first challenge, to capture the high-order correlations between nodes. Second, we propose a multi-channel cross-rating contrastive learning module, which treats different rating subgraphs as different (ratings) channels and achieves mutual enhancement by encouraging the alignment between adjacent views. Third, we also design an attention-based cross-view fusing module to further enhance the interaction between different ratings to mitigate the data sparsity and long-tail distribution challenge. Finally, we use an attention-based cross-view aggregation module to collaboratively learn node representations from local and global perspectives and fuse them to obtain the final representation. Experimental results demonstrate that MHCL achieves a +9.04% performance lift in terms of MSE compared to state-of-the-art baselines for matrix completion on ML-100K dataset. Besides, MHCL consistently outperforms advanced rating and review-based recommendation baselines on recommendation tasks, with the most notable +10.31% N@10 improvement on ML-100K and average +6.65% improvement on M@10 and N@10 across all datasets.
@@
  • MHCL introduces a multi-channel cross-rating contrastive learning approach, treating each rating subgraph as a rating channel, which strengthens the alignment between adjacent views and improves the modeling of inter-rating correlations. This approach extracts richer learning signals, improving the representations of less frequent ratings.
  • MHCL captures high-order correlations by dynamically learning hypergraph structures during training, effectively mitigating over-smoothing issues and enhancing the model’s ability to model complex correlations.
  • To mitigate data sparsity and long-tail distributions, MHCL employs an attention-based cross-view fusion module, which enhances interactions between rating subgraphs and ensures better representation of both active and inactive nodes. Moreover, we integrate contrastive learning modules to holistically capture cross-rating intrinsic correlations.
  • Extensive experiments demonstrate that MHCL significantly outperforms state-of-the-art baselines in both matrix completion and recommendation tasks, proving its effectiveness in scenarios with sparse and long-tail data distributions.
@@ 2.1. GNNs for Matrix Completion
Recently, the field of graph neural networks (GNNs) has demonstrated remarkable efficacy in the domain of matrix completion. The fundamental concept underlying GNN-based matrix completion methods is the interpretation of a matrix as a bipartite graph, where observed ratings or purchases are denoted by connecting edges, which ingeniously and effectively transforms the matrix completion problem into a predictive task centered around graph edges. Traditional GNN-based methods (Kipf and Welling, 2016; Hamilton et al., 2017; Veličković et al., 2017) acquire the information from the neighbors in the user-item graph through multi-hop convolutions to exploit the high-order relations defined by the initial pairwise links. NGCF (Wang et al., 2019) and PinSage (Ying et al., 2018) adopt multiple graph convolutions to explore high-order connectivity. LightGCN (He et al., 2020) demonstrates that neighborhood aggregation is the most essential component in GCN and proposes to omit the non-linear transformation during propagation. IDCF-NN (Wu et al., 2021b) and IDCF-GC (Wu et al., 2021b) are both inductive collaborative filtering methods; the former adopts the traditional matrix factorization method to factorize the rating matrix of a set of key users to obtain meta-latent factors, and the latter estimates the hidden relationship from query to key users and learns to use meta-latents to inductively compute the embedding of query users through neural information transfer. GC-MC (Berg et al., 2017) adopts a one-hot encoding scheme for node IDs as initial node features. It refines this representation through a GNN encoder applied to the bipartite graph and subsequently reconstructs the rating edges by employing a GNN decoder. Additionally, IMC-GAE (Shen et al., 2021) leverages identical features, role-aware attributes, and one-hot indices as initial features, enhancing the model’s generalization capabilities. In a similar vein, IGMC (Zhang and Chen, 2019) introduces a novel strategy by extracting enclosing subgraphs surrounding target edges. It employs GNNs to learn local graph patterns, enabling inductive matrix completion that generalizes to users or items not encountered during training. Taking a step further, GIMC (Zhang et al., 2022) extends inductive matrix completion into non-Euclidean spaces by incorporating hyperbolic regression. This innovation effectively mitigates the challenges posed by data sparsity. RMG (Wu et al., 2019) represents a pioneering effort, amalgamating graph signals and review information for recommender systems. Subsequently, RGCL (Shuai et al., 2022) builds upon this foundation by harnessing the semantic information embedded within review features. It fine-tunes the influence of neighboring nodes and employs contrastive learning techniques to achieve superior interaction modeling. Shuai et al. (Shuai et al., 2023) propose a novel framework TGNN, which enhances explainable recommendation by using topic modeling and GNNs to explicitly and implicitly model user preferences from reviews. This overcomes the limitations of sparsity and attribute dependency, improving both accuracy and explanation quality. MAGCL (Wang et al., 2023) uses multi-view representation learning, graph contrastive learning and multi-task learning to capture fine-grained semantic information in user reviews, and learns higher-order structure-aware and self-discriminative node embeddings.
@@
In summary, these advancements within the realm of GNNs have ushered in transformative approaches to matrix completion, showcasing their potential to revolutionize recommender systems.
2.2. Contrastive Learning
Contrastive learning has emerged as a potent self-supervised framework, and the main idea of self-supervised contrastive learning is to encourage the representation affinity among various views of the same object, while at the same time enlarging the representation dispersion of different objects. The fundamental premise of contrastive learning revolves around bringing an “anchor” and a “positive” sample closer in the embedding space while simultaneously pushing the anchor away from multiple “negative” samples (Ma et al., 2022). Drawing inspiration from the successes of contrastive learning across various domains (Chen et al., 2020; Giorgi et al., 2020; Oord et al., 2018; Cao et al., 2021; Yuan et al., 2021; Lan et al., 2019), numerous researchers have incorporated contrastive learning techniques into graph data to offer a potential solution to the challenges of insufficient supervision signals in recommender systems (Veličković et al., 2018; Zhu et al., 2020; You et al., 2020; Thakoor et al., 2021; Ren et al., 2019). SGL (Wu et al., 2021a), for instance, enriches the user-item bipartite graph with structural perturbations and subsequently focuses on maximizing the consistency of representations derived from different perspectives, as learned through a graph encoder. SGCCL (Li et al., 2023) constructs a user-user graph and an item-item graph and performs edge/feature dropout to augment them. SimGCL (Yu et al., 2022) presents experimental evidence highlighting the decisive influence of distribution consistency on recommendation performance. This study underscores the significance of distribution-based improvements over traditional dropout-based graph enhancements. It demonstrates that contrastive learning can enhance model generalization by rendering the learned representation distribution more uniform. In a parallel vein, XSimGCL (Yu et al., 2023) and RocSE (Ye et al., 2023) views by adding uniform noises to node representations, without using data augmentation. LightGCL (Cai et al., 2023) proposes a singular value decomposition-based graph augmentation strategy to effectively distill global collaborative signals.
Notably, NCL (Lin et al., 2022) delves into uncovering potential relationships among nodes by characterizing high-order potential relations as enriched neighborhoods of nodes. Tailored contrastive learning methods are devised for different types of neighborhoods, enhancing the model’s capacity to capture intricate node interactions. RGCL (Shuai et al., 2022) introduces review-based contrastive learning, selecting corresponding review representations as positive samples. This innovative strategy effectively harnesses the semantic richness encapsulated within reviews, contributing to more informed and context-aware recommendations. MAGCL (Wang et al., [n. d.]) designs a multi-aspect representation learning module to decouple multi-faceted reviews and construct multi-aspect graphs to learn high-order decoupled representations. In this way, MAGCL considers multiple semantic factors existing in reviews to represent fine-grained user preferences and item features explicitly.
In summary, contrastive learning has emerged as a promising avenue for addressing data sparsity and distribution issues within recommender systems. The application of these techniques, as demonstrated by SGL, SimGCL, NCL, and RGCL, showcases the potential of contrastive learning to augment recommendation models and enrich the user experience.
@@ 2.3. Hypergraph Learning for Recommendation
Inspired by the inherent capability of hypergraphs to transcend the limitations of fixed and heuristic graph structures in message passing (Gao et al., 2020b; Zhou et al., 2006; Bu et al., 2010; Yang et al., 2019), recent advancements in the field of recommender systems have seen the incorporation of hypergraph connections to enhance relation learning. One noteworthy model, DualHGCN (Xue et al., 2021), adeptly transforms multiplex bipartite networks into two sets of homogeneous hypergraphs. Leveraging spectral hypergraph convolutional operators, DualHGCN employs a comprehensive approach, employing intra- and inter-message passing strategies to facilitate the exchange of information within and across different domains.
Another innovative approach, MHCN (Yu et al., 2021), delves into social-based hypergraph convolutional networks for recommendations, exploiting various high-order user relations within a multi-channel framework. Additionally, DHCF (Ji et al., 2020) ventures into modeling high-order correlations among users and items for broader recommendation purposes. Nevertheless, these hypergraph-based recommendation methods necessitate the manual crafting of rules to construct heuristic hypergraph structures, which remain static throughout the training process.
In response to this limitation, researchers have begun to explore dynamic approaches for learning hypergraph structures, where global collaborative relationships are modeled during the training phase. For instance, DHSL (Zhang et al., 2018) employs labels as spatial feature vectors to aid in clustering hypergraph nodes, while DHGNN (Jiang et al., 2019) utilizes the K-NN method and K-means clustering technique to update hypergraph structures, accounting for both local and global features. Furthermore, HCCF (Xia et al., 2022) and DHLCF (Wei et al., 2022) posit that collaborative signals and topological information suffice for determining hyperedge assignments, and they employ a multi-layer perceptron (MLP) to compute these assignments. In our present work, we introduce a novel approach wherein we dynamically learn a hypergraph for each rating subgraph, enabling the capture of global collaborative relationships at a more granular level among nodes.
Table 1. The summary of the main used notations.
Notation Definition
$\mathbf{U},\mathbf{V}$ Two types of node sets (users and items)
$M$, $N$ The number of users and items
$\mathbf{R}$ The entire user-item rating matrix
$\mathcal{R}$ The set of ratings
$G$ The multiplex user-item bipartite graph
$\mathbf{E}$ The entire learnable parameter matrix
$K$ The number of hyperedges
$x^{r}$ The local embeddings of nodes within the $r$-th rating subgraph
$\mathcal{N}^{r}(i)$ The neighbors of node $i$ in $r$-th rating subgraph
$\mathbf{H}_{u}^{r},\mathbf{H}_{i}^{r}$ The adaptive hypergraph of nodes on the $r$-th rating subgraph
$h_{u}^{r},h_{i}^{r}$ The global embeddings of users and items in the $r$-th rating subgraph
$z,\Gamma$ The local and global embeddings after cross-view aggregation
$e$ The final representation of nodes
$d$ The dimension of embedding
$\mathbf{W},\mathbf{q}$ The learnable weight matrix
$\mathcal{N}_{r}$ The two ratings that are adjacent to the rating $r$